A bar of silver absorbs 1.23 kJ of energy by heat which causes its temperature to rise by 10.0°C. Calculate the specific heat of the silver bar if its mass is 525 g.



1
Expert's answer
2021-07-12T12:17:27-0400
"Q=mc\Delta T,""c=\dfrac{Q}{m\Delta T}=\dfrac{1.23\cdot10^3\ J}{0.525\ kg\cdot10^{\circ}C}=234\ \dfrac{J}{kg\cdot\!^{\circ}C}."